### v3

@@ 5 Testing Upward 1-Planarity
  • Case 1: Refer to Figure 11(a) where the thick edges schematically represent barriers $B_{1}$ (from $s_{A}$ to $s_{B}$), $B_{2}$ (from $s_{B}$ to $t_{A}$), and $B_{3}$ (from $t_{A}$ to $t_{B}$); in the figure, the red shaded shape represents $G_{A}$ and the blue shaded shape represents $G_{B}$. The figure schematically represents a series-parallel digraph $G$ with source $s_{A}$ and sink $t_{B}$. As described in Figure 11(b) an upward 1-planar drawing where the edges of $G_{A}$ precede the edges of $B_{1}$ in the left to right order of the edges exiting $s_{A}$ and the edges of $G_{B}$ precede the edges of $B_{3}$ in the left to right order of the edges entering $t_{B}$ exists if and only if $G_{A}$ crosses $G_{B}$. In fact, in any upward 1-planar drawing of $G$ that preserves its bimodal rotation system $G_{A}$ crosses $G_{B}$; conversely, starting from an upward 1-planar drawing of the union of $G_{A}$ and $G_{B}$ in which $G_{A}$ and $G_{B}$ cross, an upward 1-planar drawing of $G$ that preserves the bimodal rotation system of Figure 11(a) can be obtained as shown in Figure 11(b).
  • Case 2: We modify a very well known small digraph that is traditionally used to exhibit a digraph that is planar but not upward planar [8, 16]. As above, in Figure 11(c) thick edges schematically represent barriers, the red shaded shape represents $G_{A}$, and the blue shaded shape represents $G_{B}$. Again, an upward 1-planar drawing of this digraph exists if and only if $G_{A}$ fully crosses $G_{B}$ (see Figure 11(d)). In particular, neither $s_{B}$ nor $t_{B}$ can be drawn in the finite region of the plane bounded by $G_{A}$ since at least one $s_{A}t_{A}$-path would fully cross a barrier. Similarly, neither $s_{A}$ nor $t_{A}$ can be drawn in the finite region of the plane bounded by $G_{B}$. Observe that the digraph has a single source and a single sink and its underlying planar graph is obtained by replacing the edges of a $K_{4}$ with series-parallel graphs.
  • Case 3: Consider the digraph $G$ of Figure 11(e), where the thick edges schematically represent barriers; as in the previous case, the red shaded shape represents $G_{A}$ and the blue shaded shape represents $G_{B}$. With analogous arguments as in the previous case, it is immediate to see that $G$ has an upward 1-planar drawing if and only if $G_{A}$ fully crosses $G_{B}$ (see Figure 11(f)).
Figure 11: Some digraphs for the proof of Theorem 5.3. Thick black edges represent barriers.
The following corollary is a consequence of the argument used to prove the second case in the statement of Theorem 5.3.
Corollary 5.5.
@@
Underlying
planar graph
Acyclic orientation Upward planarity Upward 1-planarity
fixed
embedding
variable
embedding
fixed
rotation
system
variable
rotation
system
Series-parallel multi-source
multi-sink
@@
Theorem 5.3
Case 1
@@
Theorem 5.3
Case 3
single-source
single-sink
Polynomial [13, 19, 14] Trivially
polynomial
General graph multi-source
multi-sink
Polynomial
[7]
@@
[28]
[…]
Corollary 5.5
[…]
Theorem 5.3
Case 2
single-source
single-sink
Polynomial [13]
[…]
@@ 6 Testing Upward Outer-1-Planarity
Theorem 5.3 motivates the study of the complexity of testing upward 1-planarity with additional constraints. A common restriction in the study of beyond-planar graph drawing problems is the one that requires that all vertices are incident to the same face. Specifically, a graph is upward outer-1-planar if it admits an upward 1-planar embedding in which every vertex lies on the outer face. This section is devoted to the proof of the following result.
Theorem 6.1.
@@ 6.1 Basic Facts and Definitions
Conversely, let $\Gamma$ be an outer-1-planar embedding whose planarization $\Gamma^{\star}$ is acyclic. Observe that no crossing vertex of $\Gamma^{\star}$ is a source or a sink. Therefore, $\Gamma^{\star}$ only has a single source $s$ and the sinks of $G$, which are all incident to the outer face. Let $\Gamma^{+}$ be obtained from $\Gamma^{\star}$ by adding a new super sink $t$, with edges from all sinks of $G$ to $t$ into the outer face. Clearly $\Gamma^{+}$ is acyclic and a planar $st$-graph, and therefore upward planar [17]. Removing $t$ then yields an upward planar drawing of $\Gamma^{\star}$ and hence $\Gamma$ is upward outer-1-planar.
The following lemma follows from Theorem 6.2 and shows that we can decompose the problem on the biconnected components of the input graph.
Lemma 6.4.
@@
The necessity is trivial. For the sufficiency, suppose that all blocks admit an upward outer-1-planar embedding, whose planarization is thus acyclic. Combine all such planarizations at the cut-vertices so that all their vertices are incident to the outerface. Since each cycle in the graph is contained inside a block, this yields an acyclic planarization of an outer-1-planar embedding of the entire graph. This and Theorem 6.2 imply the statement.
@@
The SPQR-tree data structure, introduced by Di Battista and Tamassia [18], to represent the decomposition of a biconnected graph into its triconnected components is a special type of decomposition tree [35]. A decomposition tree of a biconnected graph $G$ is a tree $\mathcal{T}$ whose nodes $\mu$ are associated with a biconnected multi-graph $\mathrm{skel}(\mu)$, called skeleton. The edges of a skeleton can be either real or virtual, and for each node $\mu$, the virtual edges of $\mathrm{skel}(\mu)$ correspond bijectively to the edges of $\mathcal{T}$ incident to $\mu$. The tree $\mathcal{T}$ can be inductively defined as follows. In the base case, $\mathcal{T}$ consists of a single node $\mu$ whose skeleton is the graph $G$ consisting solely of real edges. In the inductive case, let $\mu$ be any node of $\mathcal{T}$ and let $H_{1}$ and $H_{2}$ be edge-disjoint connected subgraphs of $\mathrm{skel}(\mu)$ such that $\mathrm{skel}(\mu)=H_{1}\cup H_{2}$ and $H_{1}\cap H_{2}=\{u,v\}$. We obtain a new decomposition tree $\mathcal{T}^{\prime}$ from $\mathcal{T}$ by splitting the node $\mu$ into two adjacent nodes $\nu_{1}$ and $\nu_{2}$ whose skeletons are $H_{1}+uv$ and $H_{2}+uv$,respectively, where $uv$ is a new virtual edge that corresponds to the edge $\nu_{1}\nu_{2}$ of $\mathcal{T}^{\prime}$; also, we replace each edge $\mu\tau$ of $\mathcal{T}$ either with the edge $\nu_{1}\tau$ or with the edge $\nu_{2}\tau$ depending on wether edge $\mu\tau$ corresponds to a virtual edge of $H$ that belongs to $H_{1}$ or to $H_{2}$, respectively. The edges of $\mathcal{T}$ incident to $\mu$ are distributed among $\nu_{1}$ and $\nu_{2}$ based on whether their corresponding virtual edge belongs to $H_{1}$ or $H_{2}$. Let now $\mathcal{T}$ be an arbitrary decomposition tree of a biconnected graph. Note that each edge of $G$ occurs as a real edge in precisely one skeleton of $\mathcal{T}$. Consider a virtual edge $e$ in the skeleton of some node $\mu$ of $\mathcal{T}$ and let $\varepsilon$ be the edge of $\mathcal{T}$ which corresponds to $e$. The expansion graph $\mathrm{expn}(e)$ of $e$ is the graph obtained as the union of the real edges belonging to the skeletons of the subtree of $\mathcal{T}$ reachable from $\mu$ via the edge $\varepsilon$.Observe that $\mathrm{expn}(\{a,b\})$ is connected to the rest of the graph through $a$ and $b$. The vertices $a$ and $b$ are the poles of $\mathrm{expn}(\{a,b\})$. For every pair of adjacent nodes $\mu$ and $\nu$ of $\mathcal{T}$, there exists a virtual edge $e$ in $\mathrm{skel}(\mu)$ and a virtual edge $e^{\prime}$ in $\mathrm{skel}(\nu)$ with the same end-vertices. We say that $\mu$ (resp. $\nu$) is the refining node $\mathrm{refn}(e^{\prime})$ of $e^{\prime}$ (resp. $\mathrm{refn}(e)$ of $e$).
@@
Let $\mu$ be a P-node with skeleton vertices $\{u,v\}$, let $e$ be a virtual edge in $\mathrm{skel}(\mu)$ that is refined by an S-node $\lambda$ with skeleton $(u,c_{1},c_{2},\dots,c_{k},v,u)$. Then the first edge of $e$ at $u$ is the (virtual or real) edge $\{u,c_{1}\}$ of $\mathrm{skel}(\lambda)$ and the first segment of $e$ at $v$ is the (virtual or real) edge $\{c_{k},v\}$ of $\mathrm{skel}(\lambda)$.
6.2 Proof of Theorem 6.1
There are two slightly different algorithms for testing outer-1-planarity [32, 33, 4]. It is likely that both can be adapted to test upward outer-1-planarity. We choose to follow the approach by Auer et al. [4], which we briefly summarize in the following, as it seems slightly simpler to extend to our setting. Two key properties that both papers leverage are the facts that if $G$ is an outer-1-planar graph and $\mathcal{T}$ is its SPQR-tree, then the skeleton of each R-node of $\mathcal{T}$ is a $K_{4}$ that contains two non-adjacent real edges and the skeleton of each P-node has at most four virtual edges and at most one real edge. In what follows, we assume that these two properties are satisfied.
@@
By Lemma 6.4, we may assume that our input graph $G$ is biconnected and acyclic. Consider a decomposition tree $\mathcal{T}$ of $G$ (not necessarily the SPQR-tree). Let $\mu$ be a node of $\mathcal{T}$. We call an outer-1-planar embedding of $\mathrm{skel}(\mu)$ good if only real edges cross and each virtual edge is incident to the outer face. {observation} Good embeddings of all skeletons of $\mathcal{T}$ together define a unique outer-1-planar embedding of $G$, which is obtained by forming the 2-clique sums of the embeddings of the skeletons.
However, if we consider the SPQR-tree, good embeddings of the skeletons may not exist even if $G$ is outer-1-planar. Therefore, Auer et al. [4] work with a more general definition of outer-1-planar embeddings for skeletons of the SPQR-tree and show that an outer-1-planar embedding of $G$ can be combinatorially described by embeddings of the skeletons of its SPQR-tree $\mathcal{T}$ that satisfy the following five conditions C1–C5. We note that the edges of the skeletons may cross even for P-nodes, i.e., the embeddings of the skeletons we consider are not necessarily simple drawings.
@@
    Each skeleton is embedded outer-1-planar such that each virtual edge has a segment that is incident to the outer face [4, Proposition 3].
@@
    Skeletons of S-nodes are embedded planarly [4, Corollary 3].
@@
    Skeletons of R-nodes are embedded such that virtual edges are crossing-free and precisely two non-virtual edges cross each other [4, Corollary 2].
@@
    If an edge $e=\{u,v\}$ of the skeleton of a P-node is crossed, then it is a virtual edge corresponding to an S-node. Moreover, if the segment of $e$ incident to $u$ (incident to $v$) is not incident to the outer face, then the first segment of $e$ at $u$ (at $v$) must be a real edge [4, Lemma 2].
@@
    Let $\mu$ be an S-node whose skeleton contains a real edge $\{u,v\}$ and two adjacent virtual edges $\{u^{\prime},u\}$ and $\{v,v^{\prime}\}$ that are refined by $\rho$ and $\lambda$, respectively. Let $e_{\rho},e_{\lambda}$ be the virtual edges that represent $\mu$ in $\mathrm{skel}(\rho)$ and $\mathrm{skel}(\lambda)$, respectively. If both $\rho$ and $\lambda$ are P-nodes, then the segment of $e_{\rho}$ incident to $u$ must be incident to the outer face in the embedding of $\mathrm{skel}(\rho)$ or the segment of $e_{\lambda}$ incident to $v$ must be incident to the outer face in the embedding of $\mathrm{skel}(\lambda)$.
Figure 12: A portion of an SPQR-tree with an S-node $\mu$ and two adjacent S-nodes $\rho,\lambda$ illustrating the necessity of Condition 5. Since neither the segment of $e_{\rho}$ incident to $u$ is incident to the outer face in the embedding of $\mathrm{skel}(\rho)$ nor the segment of $e_{\lambda}$ incident to $v$ is incident to the outer face in the embedding of $\mathrm{skel}(\lambda)$, the real edge $\{u,v\}$ must be involved in at least two crossings. Namely, at least one crossing with an edge in the expansion graph of the blue virtual edges of $\mathrm{skel}(\rho)$ and at least one crossing with an edge in the expansion graph of the yellow virtual edges of $\mathrm{skel}(\lambda)$.
The idea behind Conditions 4 and 5 is that crossing two virtual edges $e,e^{\prime}$ in a skeleton of a P-node $\mu$ with vertices $\{u,v\}$ in such a way that $e$ is not incident to the outer face at $u$ and $e^{\prime}$ is not incident to the outer face at $v$ implies that we cross the first segment of $e$ at $u$ and the first segment of $e^{\prime}$ at $v$; see the gray-shaded regions in Figure 12 and the more formal discussion later. Condition 4 guarantees that these are real edges. On the other hand, Condition 5 ensures that each real edge receives at most one crossing in this way; see Figure 12 for an illustration of a case in which this is violated.
It follows from the work of Auer et al. [4]that together these conditions describe all outer-1-planar embeddings of $G$ without unnecessary crossings (i.e., there is no outer-1-planar drawing whose crossing edge pairs form a strict subset). We summarize this in the following theorem. As this result does not appear explicitly in the work of Auer et al., for the sake of completeness, we briefly sketch the construction. In particular, the construction of an embedding of $G$ from the embeddings of the skeletons is crucial for understanding our next steps.
Theorem 6.6.
@@
Proof 6.7.
For the bijection, observe that the necessity of 15 has been argued above. For the converse, we show how we obtain an embedding of $G$ from embeddings $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$ for each skeleton $\mathrm{skel}(\mu)$ that satisfy 15. We refer to this embedding as the combined embedding of $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$.

### v4

@@ 5 Testing Upward 1-Planarity
Figure 11: Some digraphs for the proof of Theorem 6. Thick black edges represent barriers.
  • Case 1: Refer to Fig. 11(a), where the thick edges schematically represent barriers $B_{1}$ (from $s_{A}$ to $s_{B}$), $B_{2}$ (from $s_{B}$ to $t_{A}$), and $B_{3}$ (from $t_{A}$ to $t_{B}$); in the figure, the red shaded shape represents $G_{A}$ and the blue shaded shape represents $G_{B}$. The figure schematically represents a series-parallel digraph $G$ with source $s_{A}$ and sink $t_{B}$. As described in Fig. 11(b), an upward 1-planar drawing where the edges of $G_{A}$ precede the edges of $B_{1}$ in the left to right order of the edges exiting $s_{A}$ and the edges of $G_{B}$ precede the edges of $B_{3}$ in the left to right order of the edges entering $t_{B}$ exists if and only if $G_{A}$ crosses $G_{B}$. In fact, in any upward 1-planar drawing of $G$ that preserves its bimodal rotation system $G_{A}$ crosses $G_{B}$; conversely, starting from an upward 1-planar drawing of the union of $G_{A}$ and $G_{B}$ in which $G_{A}$ and $G_{B}$ cross, an upward 1-planar drawing of $G$ that preserves the bimodal rotation system of Fig. 11(a) can be obtained as shown in Fig. 11(b).
  • Case 2: We modify a very well known small digraph that is traditionally used to exhibit a digraph that is planar but not upward planar (Bertolazzi, Di Battista, Mannino, and Tamassia, 1998; Di Battista, Eades, Tamassia, and Tollis, 1999). As above, in Fig. 11(c) thick edges schematically represent barriers, the red shaded shape represents $G_{A}$, and the blue shaded shape represents $G_{B}$. Again, an upward 1-planar drawing of this digraph exists if and only if $G_{A}$ fully crosses $G_{B}$ (see Fig. 11(d)). In particular, neither $s_{B}$ nor $t_{B}$ can be drawn in the finite region of the plane bounded by $G_{A}$ since at least one $s_{A}t_{A}$-path would fully cross a barrier. Similarly, neither $s_{A}$ nor $t_{A}$ can be drawn in the finite region of the plane bounded by $G_{B}$. Observe that the digraph has a single source and a single sink and its underlying planar graph is obtained by replacing the edges of a $K_{4}$ with series-parallel graphs.
  • Case 3: Consider the digraph $G$ of Fig. 11(e), where the thick edges schematically represent barriers; as in the previous case, the red shaded shape represents $G_{A}$ and the blue shaded shape represents $G_{B}$. With analogous arguments as in the previous case, it is immediate to see that $G$ has an upward 1-planar drawing if and only if $G_{A}$ fully crosses $G_{B}$ (see Fig. 11(f)).
The following corollary is a consequence of the argument used to prove the second case in the statement of Theorem 6.
Corollary 3.
@@
Underlying Acyclic Upward planarity Upward 1-planarity
planar orientation: Embedding Rotation system
graph sources/sinks fixed variable fixed variable
@@
single/single polynomial (Chaplick et al., 2022a, b; Didimo et al., 2010) trivially polynomial
@@
single/single polynomial (Chaplick et al., 2022a)
@@ 6 Testing Upward Outer-1-Planarity
Theorem 6 motivates the study of the complexity of testing upward 1-planarity with additional constraints. A common restriction in the study of beyond-planar graph drawing problems is the one that requires that all vertices are incident to the same face. Specifically, a graph is upward outer-1-planar if it admits an upward 1-planar embedding in which every vertex lies on the outer face. This section is devoted to the proof of the following result.
Theorem 7.
@@ 6.1 Basic Facts and Definitions
Conversely, let $\Gamma$ be an outer-1-planar embedding whose planarization $\Gamma^{\star}$ is acyclic. Observe that no crossing vertex of $\Gamma^{\star}$ is a source or a sink. Therefore, $\Gamma^{\star}$ only has a single source $s$ and the sinks of $G$, which are all incident to the outer face. Let $\Gamma^{+}$ be obtained from $\Gamma^{\star}$ by adding a new super sink $t$, with edges from all sinks of $G$ to $t$ into the outer face. Clearly $\Gamma^{+}$ is acyclic and a planar $st$-graph, and therefore upward planar (Di Battista and Tamassia, 1988). Removing $t$ then yields an upward planar drawing of $\Gamma^{\star}$ and hence $\Gamma$ is upward outer-1-planar. ∎
The following lemma follows from Theorem 8 and shows that we can decompose the problem on the biconnected components of the input graph.
Lemma 2.
@@
The necessity is trivial. For the sufficiency, suppose that all blocks admit an upward outer-1-planar embedding, whose planarization is thus acyclic. Combine all such planarizations at the cut-vertices so that all their vertices are incident to the outerface. Since each cycle in the graph is contained inside a block, this yields an acyclic planarization of an outer-1-planar embedding of the entire graph. This and Theorem 8 imply the statement. ∎
@@
The SPQR-tree data structure, introduced by Di Battista and Tamassia (1996), to represent the decomposition of a biconnected graph into its triconnected components is a special type of decomposition tree (Hopcroft and Tarjan, 1973). A decomposition tree of a biconnected graph $G$ is a tree $\mathcal{T}$ whose nodes $\mu$ are associated with a biconnected multi-graph $\operatorname{skel}(\mu)$, called skeleton. The edges of a skeleton can be either real or virtual, and for each node $\mu$, the virtual edges of $\operatorname{skel}(\mu)$ correspond bijectively to the edges of $\mathcal{T}$ incident to $\mu$. The tree $\mathcal{T}$ can be inductively defined as follows. In the base case, $\mathcal{T}$ consists of a single node $\mu$ whose skeleton is the graph $G$ consisting solely of real edges. In the inductive case, let $\mu$ be any node of $\mathcal{T}$ and let $H_{1}$ and $H_{2}$ be edge-disjoint connected subgraphs of $\operatorname{skel}(\mu)$ such that $\operatorname{skel}(\mu)=H_{1}\cup H_{2}$ and $H_{1}\cap H_{2}=\{u,v\}$. We obtain a new decomposition tree $\mathcal{T}^{\prime}$ from $\mathcal{T}$ by splitting the node $\mu$ into two adjacent nodes $\nu_{1}$ and $\nu_{2}$ whose skeletons are $H_{1}+uv$ and $H_{2}+uv$, respectively, where $uv$ is a new virtual edge that corresponds to the edge $\nu_{1}\nu_{2}$ of $\mathcal{T}^{\prime}$; also, we replace each edge $\mu\tau$ of $\mathcal{T}$ either with the edge $\nu_{1}\tau$ or with the edge $\nu_{2}\tau$ depending on wether edge $\mu\tau$ corresponds to a virtual edge of $H$ that belongs to $H_{1}$ or to $H_{2}$, respectively. The edges of $\mathcal{T}$ incident to $\mu$ are distributed among $\nu_{1}$ and $\nu_{2}$ based on whether their corresponding virtual edge belongs to $H_{1}$ or $H_{2}$. Let now $\mathcal{T}$ be an arbitrary decomposition tree of a biconnected graph. Note that each edge of $G$ occurs as a real edge in precisely one skeleton of $\mathcal{T}$. Consider a virtual edge $e$ in the skeleton of some node $\mu$ of $\mathcal{T}$ and let $\varepsilon$ be the edge of $\mathcal{T}$ which corresponds to $e$. The expansion graph $\operatorname{expn}(e)$ of $e$ is the graph obtained as the union of the real edges belonging to the skeletons of the subtree of $\mathcal{T}$ reachable from $\mu$ via the edge $\varepsilon$. Observe that $\operatorname{expn}(\{a,b\})$ is connected to the rest of the graph through $a$ and $b$. The vertices $a$ and $b$ are the poles of $\operatorname{expn}(\{a,b\})$. For every pair of adjacent nodes $\mu$ and $\nu$ of $\mathcal{T}$, there exists a virtual edge $e$ in $\operatorname{skel}(\mu)$ and a virtual edge $e^{\prime}$ in $\operatorname{skel}(\nu)$ with the same end-vertices. We say that $\mu$ (resp. $\nu$) is the refining node $\operatorname{refn}(e^{\prime})$ of $e^{\prime}$ (resp. $\operatorname{refn}(e)$ of $e$).
@@
Let $\mu$ be a P-node with skeleton vertices $\{u,v\}$, let $e$ be a virtual edge in $\operatorname{skel}(\mu)$ that is refined by an S-node $\lambda$ with skeleton $(u,c_{1},c_{2},\dots,c_{k},v,u)$. Then the first segment of $e$ at $u$ is the (virtual or real) edge $\{u,c_{1}\}$ of $\operatorname{skel}(\lambda)$ and the first segment of $e$ at $v$ is the (virtual or real) edge $\{c_{k},v\}$ of $\operatorname{skel}(\lambda)$.
6.2 Proof of Theorem 7
There exist two different algorithms for testing outer-1-planarity, namely by Hong, Eades, Katoh, Liotta, Schweitzer, and Suzuki (2013, 2015) and by Auer, Bachmaier, Brandenburg, Gleißner, Hanauer, Neuwirth, and Reislhuber (2016). It is likely that both can be adapted to test upward outer-1-planarity. We choose to follow the approach by Auer et al. (2016), which we briefly summarize in the following, as it seems slightly simpler to extend to our setting. Two key properties that both papers leverage are the facts that if $G$ is an outer-1-planar graph and $\mathcal{T}$ is its SPQR-tree, then the skeleton of each R-node of $\mathcal{T}$ is a $K_{4}$ that contains two non-adjacent real edges and the skeleton of each P-node has at most four virtual edges and at most one real edge. In what follows, we assume that these two properties are satisfied.
@@ 6.2 Proof of Theorem 7
By Lemma 2, we may assume that our input graph $G$ is biconnected and acyclic. Consider a decomposition tree $\mathcal{T}$ of $G$ (not necessarily the SPQR-tree). Let $\mu$ be a node of $\mathcal{T}$. We call an outer-1-planar embedding of $\operatorname{skel}(\mu)$ good if only real edges cross and each virtual edge is incident to the outer face.
Observation 1.
Good embeddings of all skeletons of $\mathcal{T}$ together define a unique outer-1-planar embedding of $G$, which is obtained by forming the 2-clique sums of the embeddings of the skeletons.
However, if we consider the SPQR-tree, good embeddings of the skeletons may not exist even if $G$ is outer-1-planar. Therefore, Auer et al. (2016) work with a more general definition of outer-1-planar embeddings for skeletons of the SPQR-tree and show that an outer-1-planar embedding of $G$ can be combinatorially described by embeddings of the skeletons of its SPQR-tree $\mathcal{T}$ that satisfy the following five conditions C1–C5. We note that the edges of the skeletons may cross even for P-nodes, i.e., the embeddings of the skeletons we consider are not necessarily simple drawings.
@@
    Each skeleton is embedded outer-1-planar such that each virtual edge has a segment that is incident to the outer face (Auer et al., 2016, Proposition 3).
@@
    Skeletons of S-nodes are embedded planarly (Auer et al., 2016, Corollary 3).
@@
    Skeletons of R-nodes are embedded such that virtual edges are crossing-free and precisely two non-virtual edges cross each other (Auer et al., 2016, Corollary 2).
@@
    If an edge $e=\{u,v\}$ of the skeleton of a P-node is crossed, then it is a virtual edge corresponding to an S-node. Moreover, if the segment of $e$ incident to $u$ (incident to $v$) is not incident to the outer face, then the first segment of $e$ at $u$ (at $v$) must be a real edge (Auer et al., 2016, Lemma 2).
@@
    Let $\mu$ be an S-node whose skeleton contains a real edge $\{u,v\}$ and two adjacent virtual edges $\{u^{\prime},u\}$ and $\{v,v^{\prime}\}$ that are refined by $\rho$ and $\lambda$, respectively. Let $e_{\rho},e_{\lambda}$ be the virtual edges that represent $\mu$ in $\operatorname{skel}(\rho)$ and $\operatorname{skel}(\lambda)$, respectively. If both $\rho$ and $\lambda$ are P-nodes, then the segment of $e_{\rho}$ incident to $u$ must be incident to the outer face in the embedding of $\operatorname{skel}(\rho)$ or the segment of $e_{\lambda}$ incident to $v$ must be incident to the outer face in the embedding of $\operatorname{skel}(\lambda)$.
Figure 12: A portion of an SPQR-tree with an S-node $\mu$ and two adjacent S-nodes $\rho,\lambda$ illustrating the necessity of Condition 5. Since neither the segment of $e_{\rho}$ incident to $u$ is incident to the outer face in the embedding of $\operatorname{skel}(\rho)$ nor the segment of $e_{\lambda}$ incident to $v$ is incident to the outer face in the embedding of $\operatorname{skel}(\lambda)$, the real edge $\{u,v\}$ must be involved in at least two crossings. Namely, at least one crossing with an edge in the expansion graph of the blue virtual edges of $\operatorname{skel}(\rho)$ and at least one crossing with an edge in the expansion graph of the yellow virtual edges of $\operatorname{skel}(\lambda)$.
The idea behind Conditions 4 and 5 is that crossing two virtual edges $e,e^{\prime}$ in a skeleton of a P-node $\mu$ with vertices $\{u,v\}$ in such a way that $e$ is not incident to the outer face at $u$ and $e^{\prime}$ is not incident to the outer face at $v$ implies that we cross the first segment of $e$ at $u$ and the first segment of $e^{\prime}$ at $v$; see the gray-shaded regions in Fig. 12 and the more formal discussion later. Condition 4 guarantees that these are real edges. On the other hand, Condition 5 ensures that each real edge receives at most one crossing in this way; see Fig. 12 for an illustration of a case in which this is violated.
It follows from the work of Auer et al. (2016) that together these conditions describe all outer-1-planar embeddings of $G$ without unnecessary crossings (i.e., there is no outer-1-planar drawing whose crossing edge pairs form a strict subset). We summarize this in the following theorem. As this result does not appear explicitly in the work of Auer et al., for the sake of completeness, we briefly sketch the construction. In particular, the construction of an embedding of $G$ from the embeddings of the skeletons is crucial for understanding our next steps.
Theorem 9.
@@
Proof.
For the bijection, observe that the necessity of 15 has been argued above. For the converse, we show how we obtain an embedding of $G$ from embeddings $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$ for each skeleton $\operatorname{skel}(\mu)$ that satisfy 15. We refer to this embedding as the combined embedding of $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$.